# On Picard’s Problem via Nevanlinna Theory

### v7

@@ Abstract.
We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory. That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition.
@@
Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2010 Mathematics Subject Classification:
@@ 1.1. Motivation
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values. Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]). The study of Picard’s problem on a manifold may have originated from the celebrated work due to S. T. Yau [46] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds. Yau’s work inspired people to think about the following Picard’s problem. Let be a complete non-compact Kähler manifold with non-negative Ricci curvature (see examples in Sha-Yang [35] and Tian-Yau [41, 42]). The classical Picard’s problem asks that
@@
for some such that Since is simply-connected, then the fundamental group is trivial. It implies that is necessarily a trivial group if such a lifting exists. However, is not trivial in most cases. So, we cannot lift to in general.
@@
If a complex Lie group acts on transitively, then every meromorphic function on must be a constant if it omits distinct values.
@@
It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds. Refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 16], J. Noguchi [24, 25], E. Nochka [26], M. Ru [29, 30], B. Shiffman [31], B. Shabat [32], F. Sakai [33, 34], W. Stoll [39, 40], P. Vojta [38] and H. Wu [45], and see [1, 2, 3, 10, 11, 12] and references therein.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions of slow growth rate by establishing a Nevanlinna-type theory based on heat diffusion, We shall introduce his main work. Let be the Kähler form and Laplace-Beltrami operator of associated to respectively. One can endow with the Fubini-Study metric Let be a meromorphic function. The Kählerness of implies the Hilbert-Schmidt norm of the differential (with respect to ) can be formulated as
@@
Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion. Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and utilizing Itô’s formula (see [19]), he established an analogue of Nevanlinna theory. Note that is stochastically complete (i.e., is conservative) since Grigor’yan’s criterion [14]. To make and meaningful, the following conditions are necessary:
@@
To do so, has to satisfy certain growth assumption. That is why a growth condition is needed in Theorem C.
In this paper, we generalize Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Our innovative method is to construct a family of exhaustive precompact domains for by means of a unique minimal positive global Green function for so that we can well define the Nevanlinna’s functions on and establish a Second Main Theorem by introducing the global Green function method. Apply the Second Main Theorem, we can give a positive answer to the Picard’s problem under certain condition.
@@ 1.2. Main results
Assume that is non-parabolic, by which we mean that there exists a positive global Green function for By N. Varopoulos [43, 44] (see Li-Tam-Wang [20, 22] also), the non-parabolicity implies that there exists a unique minimal positive global Green function of for denoted by where is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
holds for all and where is the Riemannian distance function of from and denotes the Riemannian volume of the geodesic ball centered at with radius in For define
Then, for all and the family exhausts
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions and on (see Section 3.2). Let denote the Ricci form of associated to the metric and be the canonical line bundle over The characteristic function of is defined by
Again, set
@@
Remark 2. Note that due to
@@
As an application of Corollary 1.2, we answer the Picard’s problem under certain condition as follows.
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
@@
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
@@ 2.1. Volume Comparison Theorem
A space form is a complete Riemannian manifold with constant sectional curvature. Let be an -dimensional (simply-connected) space form with constant sectional curvature Denote by the Riemannian volume of a geodesic ball with radius in Let be an -dimensional complete Riemannian manifold with Ricci curvature denoted by Denote by the Riemannian volume of the geodesic ball centered at a fixed point with radius in
@@
If for a constant then the volume ratio is non-increasing in and it tends to as Hence, we have
@@
holds for all where is the volume of the unit ball in
When Calabi-Yau (see [36]) proved that
@@
Assume that is non-compact. If then has an infinite volume. More precisely, for any there exists a constant such that
@@
2.2. Estimates of Heat Kernel
@@
Assume that is non-parabolic. If has non-negative Ricci curvature, then there exist constants depending only on such that
@@ 3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to the Kähler metric is defined by
@@ 3.1. Construction of
Fix a reference point Let be the Riemannian volume of where is the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
@@
By
where means we derive that is a precompact domain containing such that
In fact, the family exhausts namely, for any sequence with we have
Whence, the boundary of can be formulated as
Due to Sard’s theorem, is a submanifold of for almost all
@@
Evidently, is the positive Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
@@ 3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form where
@@
Locally, write where is a holomorphic local frame of and is a holomorphic function. By Poincaré-Lelong formula (see, e.g., [8]), it leads to
@@ 3.3. First Main Theorem
Lemma 3.1 (Jensen-Dynkin formula).
Let be a -class function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Jensen-Dynkin formula to then we are led to
Therefore, we are led to that
@@ 4. Gradient Estimates of Green Functions
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
@@
Since
we obtain
Let denote the gradient operator on a Riemannian manifold. Cheng-Yau [9] proved the following theorem.
@@
Let be a complete Riemannian manifold. Let be a non-negative harmonic function on a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
where is the Riemannian distance between and and is the lower bound of the Ricci curvature on
We have an upper bound of as follows.
@@
There exists a constant such that
@@
we can obtain an estimate for upper bounds of as follows.
@@
There exists a constant such that
where is the Riemannian area element of
@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension
@@ 5.1. Calculus Lemma
Lemma 5.1 (Borel’s Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
@@
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit
exists, here is allowed. If then Set
@@
We establish the Calculus Lemma as follows.
@@
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
@@
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant such that
@@
holds for all outside ∎
@@ 5.2. Logarithmic Derivative Lemma
in holomorphic local coordinates here is the inverse of Again, define the Nevanlinna’s characteristic function of by
@@
It is not difficult to show that
On one puts a singular metric
which leads to
@@
Note that
Whence, it concludes from Fubini’s theorem that
@@
We establish the following Logarithmic Derivative Lemma.
@@ 6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete non-compact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
where is the Ricci form of
@@
We establish a Second Main Theorem as follows.
@@
Since is of simple normal crossing type, then there are a finite open covering of and finitely many rational functions on for each such that are holomorphic on with that
@@
Fix any one can take holomorphic local coordinates near and holomorphic local coordinates near such that
Set
Then, we have and at :
@@
Moreover, put for Then, at :
Whence, at :
@@
Note that due to Corollary 2.2. We have
Using the volume growth condition, we obtain
Since a non-constant meromorphic mapping has the growth rate at least, then we have the theorem proved if is satisfied due to Theorem 6.1.
For we have
@@
So, it yields from that

### v8

@@ Abstract.
We consider the classical Picard’s problem for non-parabolic complete Kähler manifolds with non-negative Ricci curvature. Based on the global Green function approach, we give a positive answer to Picard’s problem under certain condition by developing Nevanlinna theory. That is, we prove that every meromorphic function on such a manifold reduces to a constant if it omits three distinct values, provided that the manifold satisfies a volume growth condition; and prove that every meromorphic function of non-polynomial type growth on such a manifold can avoid 2 distinct values at most.
@@
Nevanlinna theory; value distribution; Picard’s theorem; defect relation; Kähler manifolds2020 Mathematics Subject Classification:
@@ 1.1. Motivation
The famous Picard’s theorem asserts that a meromorphic function on the complex Euclidean spaces must reduce to a constant if it omits three distinct values. Historically, many authors have been committed to extensions of this theorem for a long time (see [4, 13, 15, 17, 18, 21, 28]). The study of Picard’s problem on a manifold may have originated from the celebrated work due to S. T. Yau [47] on Liouville’s theorem, which proves the Liouville’s property of harmonic functions on Ricci non-negatively curved manifolds. Yau’s work inspires people to think about the following Picard’s problem. Let be a complete noncompact Kähler manifold with non-negative Ricci curvature (see examples for such manifolds in Sha-Yang [35] and Tian-Yau [40, 41]).
The classical Picard’s problem states that
@@
for some such that Since is simply-connected, the fundamental group is trivial. This implies that is necessarily a trivial group if such a lifting admits. However, is not trivial in most cases. Thus, we cannot lift to in general.
@@
Assume that there is a complex Lie group acting on transitively. Then, every meromorphic function on must be a constant if it omits distinct values.
@@
It is known that Nevanlinna theory [27] is a great development of Picard’s theorem, which investigates the value distribution of meromorphic mappings between complex manifolds. For instance, refer to L. Ahlfors [5], H. Cartan [7], Carlson-Griffths-King [8, 16], J. Noguchi [24, 25], E. Nochka [26], M. Ru [29, 30], B. Shiffman [31], B. Shabat [32], F. Sakai [33, 34], W. Stoll [38, 39], P. Vojta [45], H. Wu [46], and also refer to [1, 2, 3, 10, 11, 12] and references therein.
In 2010, A. Atsuji [2] gave an affirmative answer for a class of meromorphic functions with slower growth by introducing a Nevanlinna-type theory based on heat diffusion, We introduce his main work. Let stand for the Kähler form and Laplace-Beltrami operator on respectively. Let be a meromorphic function. We equip with Fubini-Study metric The Kählerness of indicates that the Hilbert-Schmidt norm of differential with respect to metrics can be expressed as
@@
Atsuji’s method [2] (see [10] also) is to use stochastic calculus by Brownian motion. Atsuji successfully introduced Nevanlinna-type functions and where is the time of the Brownian motion on By computing curvatures and using Itô’s formula (see [19]), he established an analogue of the Nevanlinna theory. Note that is stochastically complete (i.e., is conservative) due to Grigor’yan’s criterion (see [14]). In order to make and meaningful, the following are necessary:
@@
Hence, needs to satisfy certain growth assumption. That is why a growth condition is added to Theorem C.
In this paper, our purpose is to study the Picard’s problem using the tool of Nevanlinna theory, i.e., we shall extend the classical Nevanlinna theory to non-parabolic complete Kähler manifolds with non-negative Ricci curvature, and apply the Second Main Theorem to establish a Picard’s theorem. Let us first recall the basic notions of the non-parabolicity and the Ricci curvature of Riemannian manifolds. Let be a complete Riemannian manifold. One says that is non-parabolic, if there exists a positive global Green function for and parabolic otherwise. Or equivalently, in a viewpoint of geometric analysis, is called non-parabolic if
for and parabolic if this infinite integral is divergent, where denotes the transition density function of the Brownian motion generated by the Laplace-Beltrami operator on and which is also called the heat kernel of Note that when is non-parabolic, defines the unique minimal positive global Green function of for More details may refer to [20, 22, 42, 43]. Let stand for the Riemannian curvature tensor of The Ricci curvature tensor of is defined by
for any with where is an orthonormal basis of The Ricci curvature of at in a direction is defined by
We say that is of non-negative Ricci curvature at if for each nonzero vector and that is of non-negative Ricci curvature, if it is of non-negative Ricci curvature at each Moreover, the sectional curvature of at along a section spanned by is defend by
From the definitions, the sign of the sectional curvature at a point determines the sign of the Ricci curvature at the same point.
To facilitate the construction of examples later for manifolds considered in this paper, we shall provide some criteria of non-parabolicity via the volume growth. A sharp necessary condition by N. Varopoulos [44] states that if is non-parabolic, then
where denotes the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in However, (1) is far from sufficient, and a counterexample was constructed in [44]. The first major result for the sufficiency was due to N. Varopoulos [43] and Li-Yau [23]. Based on the heat kernel estimates, they proved that if has non-negative Ricci curvature and (1) is satisfied, then is non-parabolic. So, for with non-negative Ricci curvature, is non-parabolic if and only if (1) is satisfied.
Next, assume that is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Here, the non-parabolicity of is meant that is non-parabolic as a Riemannian manifold. The Chern-Ricci form of on is defined by
with Since is Kähler, is equivalent to Our innovative idea in the present paper is to construct a family of exhaustive precompact domains for by means of the minimal positive global Green function for so that it allows us to well define Nevanlinna’s functions on and then establish a Second Main Theorem by estimating the gradient of Green function for Applying the Second Main Theorem, we provide a positive answer to the Picard’s problem under certain condition.
@@ 1.2. Main results
Let be the minimal positive global Green function of for in which is the Laplace-Beltrami operator on and is a fixed reference point in Using Li-Yau’s estimate [23], there are constants such that
holds for all and all in which is the Riemannian distance between and and stands for the Riemannian volume of the geodesic ball centered at with radius in For define
Then, we have for all and the family exhausts
Let be a complex projective manifold with over which one can put a Hermitian positive line bundle with Chern form We now take a divisor where denotes the complete linear system of Given a meromorphic mapping we have the Nevanlinna’s functions on (see Section 3.2). Let be the canonical line bundle over The characteristic function of is defined by
where is the Kähler form of and is the Green function of for with a pole at satisfying Dirichlet boundary condition. Set
@@
Remark 2. Note that as due to
@@
As an application of Corollary 1.2, we consider the Picard’s problem under certain condition below. Treat the case when with the point line bundle, i.e., the hyperplane line bundle Take where are distinct points in If we put then A basic and well-known fact states that
which yields that
whenever or in Corollary 1.2 is satisfied. According to the arguments above, we conclude that
@@
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. If satisfies the volume growth condition then every meromorphic function on reduces to a constant if it omits distinct values.
@@
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. Then, every meromorphic function of non-polynomial type growth on can omit distinct values at most.
To close the section, we finally provide some typical examples for where Corollaries 1.3 and 1.4 apply. According to (1), we first notice that with is a non-parabolic complete noncompact Kähler manifold with zero Ricci curvature. Below, we give two nontrivial examples.
Example 1. For an integer we equip with Fubini-Study metric Note that is a complete Kähler manifold with positive sectional curvature (and thus with positive Ricci curvature). We consider the product manifold with which is equipped with the product metric induced from and the standard Euclidean metric on Because is a complete noncompact Kähler manifold, so is Since is flat, is clearly Ricci non-negatively curved under this product metric. Again, by the dimension condition that we deduce that is of the volume growth: as which leads to that is satisfied. Therefore, is a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. It is not difficult to see that there exist a lot of nonconstant meromorphic functions on For instance, we can take or where is any meromorphic function on and is any rational function on
Example 2. Treat a product manifold with and where denotes the -dimensional complex torus equipped with the metric inherited from the standard Euclidean metric on Note that is a complete noncompact Kähler manifold with zero Ricci curvature under the induced product metric. By the similar argument as in Example 1 also shows the non-parabolicity of The existence of nonconstant meromorphic functions on can be confirmed easily. For instance, for we can take or where is any meromorphic function on and is any elliptic function with a periodic lattice on
More general, can be taken as the product of and in which is any complete compact Kähler manifold with non-negative Ricci curvature, and is a complete noncompact Kähler manifold with non-negative Ricci curvature and volume growth satisfying (1). In particular, satisfies the desired conditions.
@@ 2.1. Volume Comparison Theorem
A space form is a complete (simply-connected) Riemannian manifold with constant sectional curvature. Let denote the -dimensional space form with constant sectional curvature and let denote the Riemannian volume of a geodesic ball with radius in Let denote a complete Riemannian manifold of dimension with Ricci curvature tensor and let denote the Riemannian volume of a geodesic ball centered at a fixed reference point with radius in
@@
If then the volume ratio is a non-increasing function in and which tends to as Hence, we have
@@
holds for all where denotes the standard Euclidean volume of the unit ball in
When Calabi-Yau (see [36]) showed that
@@
Assume that is noncompact. If then has an infinite volume. More precisely, for any there exists a constant such that
@@
2.2. Estimates of Heat Kernels
@@
Assume that has non-negative Ricci curvature. If is non-parabolic, then there exist constants depending only on such that
@@ 3. Nevanlinna’s Functions and First Main Theorem
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension The Kähler form of associated to metric is defined by
@@ 3.1. Construction of
Fix a reference point Let be the Riemannian volume of where denotes the geodesic ball centered at with radius in Note that the non-parabolicity of implies that
@@
Since
one can conclude immediately that is a precompact domain containing such that whenever and that
Thus, the family exhausts i.e., for any sequence with we have
The boundary of can be described as
By Sard’s theorem, is smooth for almost all
@@
Evidently, is the positive Green function of for with a pole at satisfying Dirichlet boundary condition, i.e.,
where is the Dirac’s delta function with a pole at Let stand for the harmonic measure on with respect to defined by
where is the inward normal derivative on and is the induced Riemannian area element of
@@ 3.2. Nevanlinna’s Functions
In what follows, we will introduce Nevanlinna’s functions. Let be a meromorphic mapping, where is a complex projective manifold. Let be a Hermitian holomorphic line bundle over with the Chern form
where
so that
@@
Locally, write in which is a local holomorphic frame of and is a holomorphic function. Using Poincaré-Lelong formula (see, e.g., [8]), we get
@@ 3.3. First Main Theorem
Lemma 3.1 (Green-Dynkin formula).
Let be a function on outside a polar set of singularities at most. Assume that Then
Assume that Apply Green-Dynkin formula to we are led to
To conclude, we obtain:
@@ 4. Gradient Estimates of Green Functions
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
To give a gradient estimate of we need to introduce some lemmas as follows.
@@
Since for we obtain
Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [9] proved the following theorem.
@@
Let be a complete Riemannian manifold of dimension Let be any geodesic ball centered at with radius in Then, there exists a constant depending only on such that
holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
We obtain an upper estimate of as follows.
@@
There exists a constant independent of such that
@@
we can derive an upper estimate of as follows.
@@
There exists a constant independent of such that
where is the induced Riemannian area element of
@@ 5. Calculus Lemma and Logarithmic Derivative Lemma
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension
@@ 5.1. Calculus Lemma
Lemma 5.1 (Borel’s Growth Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that
@@
The conclusion is clearly true for Next, we assume that Since is non-decreasing, there is a number such that Using the non-decreasing property of the limit exists or If then Set
@@
We establish the following Calculus Lemma.
@@
Let be a locally integrable function on Assume that is locally bounded at Then there is a constant independent of such that for any there exists a subset of finite Lebesgue measure such that
@@
Employing Borel’s growth lemma to the left hand side of the above equality twice: one is to and another is to then we conclude that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.4 implies that there exists a constant independent of such that
@@
holds for all outside The proof is completed. ∎
@@ 5.2. Logarithmic Derivative Lemma
in a local holomorphic coordinate where is the inverse of Define the Nevanlinna’s characteristic function of by
@@
For any it is not hard to deduce that
Put a singular metric on :
@@
A direct computation gives
Whence, we conclude from Fubini’s theorem that
@@
With the above preparations, we are to prove the following Logarithmic Derivative Lemma.
@@ 6. Second Main Theorem and Defect Relation
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature, of complex dimension Let be a complex projective manifold of complex dimension not greater than Over one can put a positive holomorphic line bundle Let be a reduced divisor of simple normal crossing type. Recall that
where is the Chern-Ricci form of on
@@
We give the main theorem in this paper, i.e., the Second Main Theorem as follows.
@@
Let be a non-parabolic complete noncompact Kähler manifold with non-negative Ricci curvature. Let be a complex projective manifold of complex dimension not greater than that of Let be a reduced divisor of simple normal crossing type, where is a positive line bundle over Let be a differentiably non-degenerate meromorphic mapping. Then for any there exists a subset of finite Lebesgue measure such that
@@
Since has the simple normal crossing type, there exist a finite open covering of and finitely many rational functions on for each such that are holomorphic on with
@@
Fix any Take a local holomorphic coordinate near and a local holomorphic coordinate near such that
and
Set Then, we have and
@@
Moreover, put for Then
which yields that
@@
We first show that the corollary holds if is satisfied. By Corollary 2.2, we have Thus, we are led to that
Using the condition we obtain
Since each non-constant meromorphic mapping has growth at least, we have the theorem proved if is satisfied due to Theorem 6.1. For we obtain
@@
Thus, it yields from that